OpenStudy (anonymous):

x^-9 * x^-7 write your answer with a positive exponent only

OpenStudy (anonymous):

@gladys2519

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (campbell_st):

well multiplying add the powers \[x^a \times x^b = x^{a + b}\] use this index law... hope it helps

OpenStudy (anonymous):

@TrapBoiTrynaEducaate

OpenStudy (anonymous):

16

OpenStudy (anonymous):

x^16

OpenStudy (anonymous):

@campbell_st so x^16

OpenStudy (campbell_st):

almost... since both powers are negative it will be -16

OpenStudy (anonymous):

But its says to write your answer with a positive exponent only

OpenStudy (campbell_st):

yes... so the negative indicates a fraction \[x^{-a} = \frac{1}{x^a}\] now use this index law

OpenStudy (campbell_st):

just replace a with 16

OpenStudy (anonymous):

So 1/x^16 is the answer

OpenStudy (campbell_st):

that's correct